# Surface topological quantum criticality: Conformal manifolds and Discrete Strong Coupling Fixed Points

### v1

@@ IX Conclusion
B) The infrared stable conformal manifolds In $D_{p}-1$dimension phase boundaries, there can be multiple conformal manifolds, as well as isolated strong coupling fixed points with very different infrared stabilities. The infrared stable conformal manifolds mentioned in A), which play the paramount role in our studies of surface topological quantum critical points, are always uniquely characterized by a finite number of marginal operators $n_{MAR}$. The dimension of these smooth manifolds itself is precisely defined by $n_{Mar}$, the number of marginal operators associated with the manifold.
@@
It is worth mentioning that there appears to be an intimate connection between the geometry of the conformal manifolds in the parameter space and the emergent dynamic symmetries of the RG equation. Assume the symmetry group of the RG equation is $G$ and the invariant subgroup of a fixed point is H, we have found that the corresponding conformal manifold is precisely the coset space of G induced by the subgroup H, i.e. $G/H$ (see Section.V). However, this emergent dynamic symmetry group G is sufficient but not necessary for the appearance of the conformal manifolds. In our studies, we also show that conformal manifolds are robust to the effects of certain class of G-symmetry breaking terms, although there can be various distortions.
@@
E) Emergent boson fields at conformal manifolds At the infrared stable conformal manifolds, we show that the four-fermion interactions are factorizable in the $\mathcal{N}$-dimension flavour space. Therefore, these conformal manifolds at the phase boundaries can be more conveniently studied via an emergent boson field that interacts with gapless surface fermions (see Eqns.49,80).
@@ A.1 Domain wall model
In this section, we develop an effective low-energy continuum model for the surface states. We show the derivation explicitly for a specific case of an isotropic cubic lattic with a single Dirac cone surface and then argue that the results can be extended to more general cases of TI surfaces with $\mathcal{N}$ flavors of surface fermions.
Consider an isotropic case of the bulk Hamiltonian given in Eqn.1, which is when $t_{x}=t_{y}=t_{z}=t$. In this case, when $11/3<t<$ 1, the TI is in a non-trivial phase and the surface normal to $z$-axis hosts a Dirac cone at the TRI point $\textbf{Q}_{1}=(0,0)$. In the bulk BZ, the mass gap is negative at the TRI point $\textbf{K}=(0,0,0)$, while it is positive at all the remaining seven bulk TRI points. In this context, to derive the surface states, we shall expand the bulk Hamiltonian to first order in k about K to get,
@@
where we define $\Gamma_{s}=\beta\alpha_{z}$ and $\varphi_{1}(k_{x},k_{y},z)$ is a 4-component spinor wavefunction. This implies that the surface state if it exists at $z=0$ must be an eigenstate of the operator $\Gamma_{s}$ with eigenvalue $+1$. That is, $\Gamma_{s}\varphi_{1}=+\varphi_{1}$. Now projecting the Hamiltonian to the two-dimensional space formed by the surface states having positive eigenvalue to the $\Gamma_{s}$ operator, we arrive at the following effective surface Hamiltonian,
@@
The eigenstates will be 2-component spinors. This is exclusively for the low-energy surface fermions in and around the surface TRI point $\textbf{Q}_{1}=(0,0)$. The effective Hamiltonians of the surface fermions at the other three TRI points can be derived similarly. The corresponding Lagrangians are given in Eqn.21 in the main text.
@@
On comparing this with Eqn.91, we see that the surface state, if it exists, must be an eigenstate of the $\Gamma_{s}$ operator but with eigenvalue $-1$. The effective surface Hamiltonian will still have the same form as in Eqn.91, but the difference here is that the pseudo-chirality of the surface state is different.
@@ A.2 Effective 2D interacting theory
where $f_{i}(z)=\sqrt{m_{i}}e^{-m_{i}|z|}$. This is obtained by solving the Schrodinger equation given in Eqn.91 for the surface states. Integrating out $z$, we arrive at an effective 2-dimensional theory of attractively interacting surface fermions as,
@@
Appendix B Renormalization group study of the interacting TI surface in $D=2+\epsilon$ dimensions up to two-loop level
Here, we show the derivation of the RG equation for a TI surface with generic $\mathcal{N}$ flavors of attractively interacting surface fermions. First, we identify the Lagrangian shown in Eqns.21,LABEL:Lfrfmn as the bare Lagrangian with bare fields $\psi_{i,0}$’s and bare couplings $V^{0}_{ij}$. Here we write down the renormalized Lagrangian as follows,
@@
To facilitate discussions on the RGE, we make a few definitions here. From Eqn.99, we can see that the $bare$ dimensionless interaction matrix element has the following form,
@@
Plugging this back to the RGE in Eqn.99 taking derivative w.r.t $\ln\mu$, we have
@@
$\sum_{l,m}\beta_{lm}\frac{\partial\tilde{V}^{0}_{ij}}{\partial\tilde{V}^{R}_{ml}}$ $=$ $\epsilon\tilde{V}_{ij}$ (103)
@@ B.1 Vertex renormalization
Here $\textbf{q}=\textbf{k}_{1}+\textbf{k}_{2}=\textbf{k}_{3}+\textbf{k}_{4}$, where $\textbf{k}_{1},\textbf{k}_{2}$ and $\textbf{k}_{1},\textbf{k}_{4}$ are the incoming and outgoing momenta respectively in the $2+1D$ spacetime. We evaluate the integral using dimensional regularization technique at arbitrary $D$-spacetime dimensions. Then, we write $D=2+\epsilon$ and expand the function to zeroth order in $\epsilon$. We get,
$\Gamma_{ij}^{(4)}(\textbf{q},\tilde{V}_{ij})=$ (106)
$-\mu^{-\epsilon}\tilde{V}_{ij}-\frac{4}{\pi}\mu^{-2\epsilon}\sum_{l}\tilde{V}_{il}\tilde{V}_{lj}\left[-\frac{2}{\epsilon}+\ln\frac{1}{\textbf{q}^{2}}-\gamma+1+\mathcal{O}(\epsilon)\right]$
@@
Now we use the relation between the bare and the renormalized coupling constants derived in Eqn.100 to write down the renormalized four-point function in terms of the renormalized interaction strength. Demanding the renormalized four-point function to be finite up to second order in the interaction strength, the vertex renormalization factor $Z^{V}_{ij}$ becomes,
@@
B.2 Field strength renormalization
@@
where $C$ is a numerical constant. Plugging this back to Eqn.109, we get,
@@
B.3 Renormalization group equation
Plugging these results back to Eqn.100, we obtain the bare coupling constant $\tilde{V}^{0}_{ij}$ as a function of the renormalized coupling constant $\tilde{V}^{R}_{ij}$. $\tilde{V}^{0}_{ij}$ is then expanded in powers of $\tilde{V}^{R}_{ij}$ to the third order. Then, we take the derivative with respect to the $\tilde{V}^{R}_{ij}$ and plug the results into Eqn.103 to obtain the beta function as,
@@ B.3 Renormalization group equation
Having derived the RGE, we remark on the definitions/conventions used in the main text for the Eqn.115 to simplify the discussions. In the main text, we use $V_{ij}$ to represent the dimensionless renormalized coupling instead of $\tilde{V}^{R}_{ij}$. We rescaled the coupling matrix as $\frac{16}{2\pi}V_{ij}\rightarrow V_{ij}$. Also, the flow parameter $\mu$ is redefined as $\mu=e^{-l}$. With these definitions, the RGE of Eqn.115 takes the form,
@@
In the $\mathcal{N}=2$ case, the RGE is the same as shown in Eqn.88 So we find that the parameter $a$ defined in Eqn.88 has the value $a=1/32$. If the two-loop field renormalization effects are ignored, then the above RGE is reduced to Eqn.27.
Using Eqn,104, we calculate the anomalous dimension of the surface fermion of flavor $i$ as,
@@
Appendix C Table of scaling dimensions of the operators at the selected points in the conformal manifolds
Below we shall list out the set of relevant, irrelevant, and marginal operators and their respective scaling dimensions at some selected points on the conformal manifold. For brevity, we denote the Cooper pair operator by the form
@@ Appendix C Table of scaling dimensions of the operators at the selected points in the conformal manifolds
We also define $\delta\hat{V}=\hat{V}-\hat{V}_{c}$.
C.1 $\mathcal{N}=2$ TI surface
Table 6: Irrelevant, marginal and relevant operators at a point $(V^{c}_{0},V^{c}_{z},V^{c}_{x})=(\frac{\epsilon}{2},\frac{\epsilon}{2}\cos\phi,\frac{\epsilon}{2}\sin\phi)$ on the ring manifold.
@@ C.1 $\mathcal{N}=2$ TI surface
Irrelevant operators $\left(\delta V_{z}\sin\phi-\delta V_{0}\right)\Phi^{\dagger}_{1}\Phi_{1}+\delta V_{x}\cos\phi\,\Phi^{\dagger}_{1}\Phi_{2}$ $-\left(\delta V_{z}\sin\phi+\delta V_{0}\right)\Phi^{\dagger}_{2}\Phi_{2}$ + h.c $-\epsilon$
Marginal operators $\delta V_{z}\cos\phi\Phi^{\dagger}_{1}\Phi_{1}-\delta V_{x}\sin\phi\Phi^{\dagger}_{1}\Phi_{2}$ $-\delta V_{z}\cos\phi\Phi^{\dagger}_{2}\Phi_{2}$ + h.c 0
Relevant operators $\left(\delta V_{z}\sin\phi+\delta V_{0}\right)\Phi^{\dagger}_{1}\Phi_{1}+\delta V_{x}\cos\phi\Phi^{\dagger}_{1}\Phi_{2}$ $\left(-\delta V_{z}\sin\phi+\delta V_{0}\right)\Phi^{\dagger}_{2}\Phi_{2}$ + h.c $\epsilon$
@@ C.2 $\mathcal{N}=3$ TI surface
Table 7: Relevant, irrelevant, marginal operators and their respective scaling dimensions at $\hat{V}_{c}=\left(\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&\frac{\epsilon}{2}\end{array}\right)$( which is the $\theta=\phi=0$ point on the manifold ).
@@
Irrelevant operators $\delta V_{22}\Phi^{\dagger}_{2}\Phi_{2}$ $-\epsilon$
@@
Relevant operators $\delta V_{33}\Phi^{\dagger}_{3}\Phi_{3}$ $\epsilon$
Table 8: Relevant, irrelevant, marginal operators and their respective scaling dimensions at $\hat{V}_{c}=\frac{\epsilon}{2}\left(\begin{array}[]{ccc}1&1&0\\ 1&1&0\\ 0&0&0\end{array}\right)$ ( which is the $\theta=\frac{\pi}{2},\phi=\frac{\pi}{4}$ point on the manifold).
@@
Irrelevant operators $\delta V_{33}\Phi^{\dagger}_{3}\Phi_{3}$ $-\epsilon$
@@
$\delta V_{11}\Phi^{\dagger}_{1}\Phi_{1}$ $-\delta V_{12}\Phi^{\dagger}_{1}\Phi_{2}$ + $\delta V_{22}\Phi^{\dagger}_{2}\Phi_{2}$ + h.c $-\epsilon$
@@
Table 9: Relevant, irrelevant, marginal operators and their respective scaling dimensions at $\hat{V}_{c}=\left(\begin{array}[]{ccc}\epsilon&0&0\\ 0&\epsilon&0\\ 0&0&0\end{array}\right)$ ( which is the $\theta=\phi=0$ point on the manifold).
@@
Irrelevant operators $\delta V_{33}\Phi^{\dagger}_{3}\Phi_{3}$ $-\epsilon$
@@
Relevant operators $\delta V_{11}\Phi^{\dagger}_{1}\Phi_{1}$ $\epsilon$
@@
$\delta V_{22}\Phi^{\dagger}_{2}\Phi_{2}$ $\epsilon$
Table 10: Relevant, irrelevant, marginal operators and their respective scaling dimensions at $\hat{V}_{c}=\frac{\epsilon}{2}\left(\begin{array}[]{ccc}1&-1&0\\ -1&1&0\\ 0&0&2\end{array}\right)$ (which is the $\theta=\frac{\pi}{2},\phi=\frac{\pi}{4}$ point on the manifold).
@@
$-\delta V_{11}\Phi^{\dagger}_{1}\Phi_{1}+\delta V_{22}\Phi^{\dagger}_{2}\Phi_{2}$ 0
Relevant operators $\delta V_{33}\Phi^{\dagger}_{3}\Phi_{3}$ $\epsilon$
@@
$\delta V_{11}\Phi^{\dagger}_{1}\Phi_{1}-\delta V_{12}\Phi^{\dagger}_{1}\Phi_{2}+\delta V_{22}\Phi^{\dagger}_{2}\Phi_{2}$ $\epsilon$
Appendix D Table of scaling dimensions at the isolated fixed points after symmetry-breaking two-loop effects are included
In section VIIIB, we found that the ring conformal manifold breaks down into isolated fixed points when two-loop fermion field renormalization effects are included. In Tables.11,12, we list the scaling dimensions of the relevant and irrelevant operator(s) at two of these four fixed points. Notice the presence of an operator at the fixed point with its scaling dimension proportional to $a\epsilon^{2}$, second order in $\epsilon$. It suggests that the operator was marginal at the one-loop level.
Table 11: Scaling dimensions of the relevant and irrelevant operators at the isolated fixed point $\hat{V}_{c}=\left(\frac{\epsilon}{2}+a\epsilon^{2}\right)\left(\hat{I}+\sigma_{z}\right)$
@@
Irrelevant operators $\left(\delta V_{z}-\delta V_{0}\right)\Phi^{\dagger}_{1}\Phi_{1}-\left(\delta V_{z}+\delta V_{0}\right)\Phi^{\dagger}_{2}\Phi_{2}$ $-\epsilon$
Relevant operators $\left(\delta V_{0}+\delta V_{z}\right)\Phi^{\dagger}_{1}\Phi_{1}+\left(\delta V_{0}-\delta V_{z}\right)\Phi^{\dagger}_{2}\Phi_{2}$ $\epsilon-2a\epsilon^{2}$
$\delta V_{x}\Phi^{\dagger}_{1}\Phi_{2}$+ h.c $a\epsilon^{2}$
Table 12: Scaling dimensions of the relevant and irrelevant operators at the isolated fixed point $\hat{V}_{c}=\left(\frac{\epsilon}{2}+\frac{a}{2}\epsilon^{2}\right)\left(\hat{I}+\sigma_{x}\right)$.
@@
Irrelevant operators $-\delta V_{0}\left(\Phi^{\dagger}_{1}\Phi_{1}+\Phi^{\dagger}_{2}\Phi_{2}\right)+\delta V_{x}\Phi^{\dagger}_{1}\Phi_{2}$ + h.c $-\epsilon-a\epsilon^{2}$
$\delta V_{z}\left(\Phi^{\dagger}_{1}\Phi_{1}-\Phi^{\dagger}_{2}\Phi_{2}\right)$ $-a\epsilon^{2}$
Relevant operators $\delta V_{0}\left(\Phi^{\dagger}_{1}\Phi_{1}+\Phi^{\dagger}_{2}\Phi_{2}\right)+\delta V_{x}\Phi^{\dagger}_{1}\Phi_{2}$ + h.c $\epsilon-a\epsilon^{2}$

### v3

@@ X Conclusion
B) The infrared stable conformal manifolds In $D_{p}-1$dimension phase boundaries, there can be multiple conformal manifolds, as well as isolated strong coupling fixed points with very different infrared stabilities. The infrared stable conformal manifolds mentioned in A), which play a paramount role in our studies of surface topological quantum critical points, are always uniquely characterized by a finite number of marginal operators $n_{MAR}$. The dimension of these smooth manifolds itself is precisely defined by $n_{MAR}$, the number of marginal operators associated with the manifold.
One important property of a marginal operator $\mathcal{O}_{M}(x)$ is that its scaling dimension $\triangle_{\mathcal{O}_{M}}=D$, the spacetime dimensions of the theory. Hence, the correlation function of the marginal operators at their respective conformal manifold obeys the power law $<\mathcal{O}_{M}(x)\mathcal{O}(0)>\sim\frac{1}{|x|^{2D}}$. Verifying this scaling behavior experimentally could provide direct evidence for the existence of conformal manifolds.
@@
It is worth mentioning that there appears to be an intimate connection between the geometry of the conformal manifolds in the parameter space and the emergent dynamic symmetries of the RG equation. Assume the symmetry group of the RG equation is $G$ and the invariant subgroup of a fixed point is H, we have found that the corresponding conformal manifold is precisely the coset space of G induced by the subgroup H, i.e. $G/H$ (see Section.VI). However, this emergent dynamic symmetry group G is sufficient but not necessary for the appearance of the conformal manifolds. In our studies, we also show that conformal manifolds are robust to the effects of certain class of G-symmetry breaking terms, although there can be various distortions.
@@
E) Emergent boson fields at conformal manifolds At the infrared stable conformal manifolds, we show that the four-fermion interactions are factorizable in the $\mathcal{N}$-dimension flavour space. Therefore, these conformal manifolds at the phase boundaries can be more conveniently studied via an emergent boson field that interacts with gapless surface fermions (see Eqns.52,84).
@@ A.1 Domain wall model
In this section, we develop an effective low-energy continuum model for the surface states. We show the derivation explicitly for a specific case of an isotropic cubic lattice with a single Dirac cone surface and then argue that the results can be extended to more general cases of TI surfaces with $\mathcal{N}$ flavors of surface fermions.
Consider an isotropic case of the bulk Hamiltonian given in Eqn.2, which is when $t_{x}=t_{y}=t_{z}=t$. In this case, when $11/3<t<$ 1, the TI is in a non-trivial phase and the surface normal to $z$-axis hosts a Dirac cone at the TRI point $\textbf{Q}_{1}=(0,0)$. In the bulk BZ, the mass gap is negative at the TRI point $\textbf{K}=(0,0,0)$, while it is positive at all the remaining seven bulk TRI points. In this context, to derive the surface states, we shall expand the bulk Hamiltonian to first order in k about K to get,
@@
where we define $\Gamma_{s}=\beta\alpha_{z}$ and $\varphi_{1}(k_{x},k_{y},z)$ is a 4-component spinor wavefunction. This implies that the surface state, if it exists at $z=0$, must be an eigenstate of the operator $\Gamma_{s}$ with eigenvalue $+1$. That is, $\Gamma_{s}\varphi_{1}=+\varphi_{1}$. Now, projecting the Hamiltonian to the two-dimensional subspace formed by the surface states having positive eigenvalue to the $\Gamma_{s}$ operator, we arrive at the following effective surface Hamiltonian,
@@
The eigenstates will be 2-component spinors. This is exclusively for the low-energy surface fermions in and around the surface TRI point $\textbf{Q}_{1}=(0,0)$. The effective Hamiltonians of the surface fermions at the other three TRI points can be derived similarly. The corresponding Lagrangians are given in Eqn.22 in the main text.
@@
On comparing this with Eqn.95, we see that the surface state, if it exists, must be an eigenstate of the $\Gamma_{s}$ operator but with eigenvalue $-1$. The effective surface Hamiltonian will still have the same form as in Eqn.95, but the difference here is that the pseudo-chirality of the surface state is different.
@@ A.2 Effective 2D interacting theory
where $f_{i}(z)=\sqrt{m_{i}}e^{-m_{i}|z|}$. This is obtained by solving the Schrodinger equation given in Eqn.95 for the surface states. Integrating out $z$, we arrive at an effective 2-dimensional theory of attractively interacting surface fermions as,
@@
Figure 12: The figure illustrates how the unitary operator $U_{1}(k_{z},\lambda)$ (Eqn.18a) changes the bulk topology. a) Mass signs ($\delta_{\Gamma_{i}}$) at the eight TRI points in the bulk BZ before the unitary operation. The topological phase is $(1;1,1,1)$. b) Bulk mass signs after the unitary operation. The bulk topology changes to $(1;0,0,1)$. c) and d) show the surface spectrum found by solving the Hamiltonian in Eqn.102 before ($(\lambda=0)$) and after ($\lambda=1$) the unitary evolution, respectively. $\textbf{Q}_{i}$’s are surface TRI points (See Fig.1 for exact definition).
Appendix B Action of non-local unitary operator $U_{1}(k_{z},\lambda)$ on the TI bulk
Here, we study the action of the unitary operator $U_{1}(k_{z},\lambda)$ (Eqn.18a) on the bulk Hamiltonian in Eqn.2. In the main text, we argued that the unitary operation changes the bulk topology in addition to deforming the surface states. We show this explicitly here by using a practical example (See Fig.12).
Consider the isotropic case of the Hamiltonian in Eqn.2, just as in the main text. That is, we set $t_{i}=t\,\forall i=x,y,z$. Under the unitary operation, the Hamiltonian evolves as,
^$\hat{\mathcal{H}}_{\textbf{K}}(\textbf{k},\lambda)$ $=$ $U_{1}^{-1}(k_{z},\lambda)\hat{\mathcal{H}}_{\textbf{K}}(\textbf{k})U_{1}(k_{z},\lambda)$ (102)
$=$ $\sigma_{z}\left(\alpha_{x}\sin k_{x}+\alpha_{y}\sin k_{y}\right)\left(\cos\lambda\pi\phi_{\textbf{k}}+i\sigma_{y}\otimes\gamma\sin\lambda\pi\phi_{\textbf{k}}\right)+\alpha_{z}\sin k_{z}-\sigma_{z}\beta(t\cos k_{x}+t\cos k_{y})$
$+$ $\beta(1-t\cos k_{z})\left(\cos\lambda\pi\phi_{\textbf{k}}+i\sigma_{y}\otimes\gamma\sin\lambda\pi\phi_{\textbf{k}}\right)$
We observe that the unitary operation is non-local and breaks both translation and inversion symmetries for intermediate values of $\lambda$. While both the symmetries are restored when $\lambda$ approaches unity, the Hamiltonian still has infinite-range hopping terms. In short, the unitary operation possesses the same set of properties as $U(k_{z},\lambda)$ (Eqn.18c), the one discussed in the main text (see Section.IVC).
Fig.12 illustrates the action of the unitary operator on the TI bulk characterized by the topological phase $(1;1,1,1)$. We see that the mass signs ($\delta_{\Gamma_{i}}$’s) (Eqn.3a) at the $k_{z}=\pi$ plane in the bulk BZ flips their sign. However, it is simultaneously followed by an interchange of the mass points separated by $\textbf{K}=(\pi,\pi,0)$, resulting in the change of bulk topological invariants. The change of the topological invariants is not surprising as the three weak invariants $(v_{1}v_{2}v_{3})$ are robust only if the crystal translation symmetry is respected. In this case, we break the crystal translation symmetry by introducing momentum coupling, and the unitary operation is non-local.
Figs.12c,12d show the surface spectrum before and after the unitary evolution, respectively. The unitary operation deforms the TI surface from a single Dirac cone phase to three Dirac cones phase. However, unlike the full unitary operation discussed in the main text, the location of the Dirac cones is not complementary to each other on the surface BZ.
Appendix C Renormalization group study of the interacting TI surface in $D=2+\epsilon$ dimensions up to two-loop level
Here, we show the derivation of the RG equation for a TI surface with generic $\mathcal{N}$ flavors of attractively interacting surface fermions. First, we identify the Lagrangian shown in Eqns.22,LABEL:Lfrfmn as the bare Lagrangian with bare fields $\psi_{i,0}$’s and bare couplings $V^{0}_{ij}$. Here we write down the renormalized Lagrangian as follows,
@@ Appendix B Action of non-local unitary operator $U_{1}(k_{z},\lambda)$ on the TI bulk
To facilitate discussions on the RGE, we make a few definitions here. From Eqn.104, we can see that the $bare$ dimensionless interaction matrix element has the following form,
@@
Plugging this back to the RGE in Eqn.104 taking derivative w.r.t $\ln\mu$, we have
@@
$\sum_{l,m}\beta_{lm}\frac{\partial\tilde{V}^{0}_{ij}}{\partial\tilde{V}^{R}_{ml}}$ $=$ $\epsilon\tilde{V}^{0}_{ij}$ (108)
@@ C.1 Vertex renormalization
Here $\textbf{q}=\textbf{k}_{1}+\textbf{k}_{2}=\textbf{k}_{3}+\textbf{k}_{4}$, where $\textbf{k}_{1},\textbf{k}_{2}$ and $\textbf{k}_{1},\textbf{k}_{4}$ are the incoming and outgoing momenta respectively in the $2+1D$ spacetime. We evaluate the integral using the dimensional regularization technique at arbitrary $D$-spacetime dimensions. Then, we write $D=2+\epsilon$ and expand the function to zeroth order in $\epsilon$. We get,
$\Gamma_{ij}^{(4)}(\textbf{q},\tilde{V}^{0}_{ij})=$ (111)
$-\mu^{-\epsilon}\tilde{V}^{0}_{ij}-\frac{4}{\pi}\mu^{-2\epsilon}\sum_{l}\tilde{V}^{0}_{il}\tilde{V}^{0}_{lj}\left[-\frac{2}{\epsilon}+\ln\frac{1}{\textbf{q}^{2}}-\gamma+1+\mathcal{O}(\epsilon)\right]$
@@
Now we use the relation between the bare and the renormalized coupling constants derived in Eqn.105 to write down the renormalized four-point function in terms of the renormalized interaction strength. Demanding the renormalized four-point function to be finite up to second order in the interaction strength, the vertex renormalization factor $Z^{V}_{ij}$ becomes,
@@
C.2 Field strength renormalization
@@
where $C$ is a numerical constant. Plugging this back to Eqn.114, we get,
@@
C.3 Renormalization group equation
Plugging these results back to Eqn.105, we obtain the bare coupling constant $\tilde{V}^{0}_{ij}$ as a function of the renormalized coupling constant $\tilde{V}^{R}_{ij}$. $\tilde{V}^{0}_{ij}$ is then expanded in powers of $\tilde{V}^{R}_{ij}$ to the third order. Then, we take the derivative with respect to the $\tilde{V}^{R}_{ij}$ and plug the results into Eqn.108 to obtain the beta function as,
@@ C.3 Renormalization group equation
Having derived the RGE, we remark on the definitions/conventions used in the main text for the Eqn.120 to simplify the discussions. In the main text, we use $V_{ij}$ to represent the dimensionless renormalized coupling instead of $\tilde{V}^{R}_{ij}$. We rescaled the coupling matrix as $\frac{16}{2\pi}V_{ij}\rightarrow V_{ij}$. Also, the flow parameter $\mu$ is redefined as $\mu=e^{-l}$. With these definitions, the RGE of Eqn.120 takes the form,
@@
In the $\mathcal{N}=2$ case, the RGE is the same as shown in Eqn.92 So we find that the parameter $a$ defined in Eqn.92 has the value $a=1/32$. If the two-loop field renormalization effects are ignored, then the above RGE is reduced to Eqn.28.
Using Eqn,109, we calculate the anomalous dimension of the surface fermion of flavor $i$ as,
@@
Appendix D Table of scaling dimensions of the operators at the selected points in the conformal manifolds
Below, we shall list the set of relevant, irrelevant, and marginal operators and their respective scaling dimensions at some selected points on the conformal manifold. For brevity, we denote the Cooper pair operator by the form
@@ Appendix D Table of scaling dimensions of the operators at the selected points in the conformal manifolds
D.1 $\mathcal{N}=2$ TI surface
Table 6: Irrelevant, marginal and relevant operators at a point $(V^{c}_{0},V^{c}_{z},V^{c}_{x})=(\frac{\epsilon}{2},\frac{\epsilon}{2}\cos\phi,\frac{\epsilon}{2}\sin\phi)$ on the ring manifold. Here, $D$ is the spacetime dimensions.
@@ D.1 $\mathcal{N}=2$ TI surface
Irrelevant operators $\left(\cos\phi-1\right)\Phi^{\dagger}_{1}\Phi_{1}+\sin\phi\,\Phi^{\dagger}_{1}\Phi_{2}$ $-\left(\cos\phi+1\right)\Phi^{\dagger}_{2}\Phi_{2}$ + h.c $22D-$ 2
Marginal operators $\cos\phi\Phi^{\dagger}_{1}\Phi_{1}-\sin\phi\Phi^{\dagger}_{1}\Phi_{2}$ $-\cos\phi\Phi^{\dagger}_{2}\Phi_{2}$ + h.c $D$
Relevant operators $\left(1+\cos\phi\right)\Phi^{\dagger}_{1}\Phi_{1}+\sin\phi\Phi^{\dagger}_{1}\Phi_{2}$ $\left(1-\cos\phi\right)\Phi^{\dagger}_{2}\Phi_{2}$ + h.c $2$
@@ D.2 $\mathcal{N}=3$ TI surface
Table 7: Relevant, irrelevant, marginal operators and their respective scaling dimensions at $\hat{V}_{c}=\left(\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&\epsilon\end{array}\right)$($\theta=\phi=0$ point on the manifold ). Here, $D$ is the spacetime dimensions.
@@
Irrelevant operators $\Phi^{\dagger}_{2}\Phi_{2}$ $22D-$ 2
@@
Relevant operators $\Phi^{\dagger}_{3}\Phi_{3}$ $2$
Table 8: Relevant, irrelevant, marginal operators and their respective scaling dimensions at $\hat{V}_{c}=\frac{\epsilon}{2}\left(\begin{array}[]{ccc}1&1&0\\ 1&1&0\\ 0&0&0\end{array}\right)$ ( $\theta=\frac{\pi}{2},\phi=\frac{\pi}{4}$ point on the manifold). Here, $D$ is the spacetime dimensions.
@@
Irrelevant operators $\Phi^{\dagger}_{3}\Phi_{3}$ $22D-$ 2
@@
$\Phi^{\dagger}_{1}\Phi_{1}$ $-\Phi^{\dagger}_{1}\Phi_{2}$ + $\Phi^{\dagger}_{2}\Phi_{2}$ + h.c $22D-$ 2
@@
Table 9: Relevant, irrelevant, marginal operators and their respective scaling dimensions at $\hat{V}_{c}=\left(\begin{array}[]{ccc}\epsilon&0&0\\ 0&\epsilon&0\\ 0&0&0\end{array}\right)$ ( $\theta=\phi=0$ point on the manifold). Here, $D$ is the spacetime dimensions.
@@
Irrelevant operators $\Phi^{\dagger}_{3}\Phi_{3}$ $22D-$ 2
@@
Relevant operators $\Phi^{\dagger}_{1}\Phi_{1}$ $2$
@@
$\Phi^{\dagger}_{2}\Phi_{2}$ $2$
Table 10: Relevant, irrelevant, marginal operators and their respective scaling dimensions at $\hat{V}_{c}=\frac{\epsilon}{2}\left(\begin{array}[]{ccc}1&-1&0\\ -1&1&0\\ 0&0&2\end{array}\right)$ ( $\theta=\frac{\pi}{2},\phi=\frac{\pi}{4}$ point on the manifold). Here, $D$ is the spacetime dimensions.
@@
$-\Phi^{\dagger}_{1}\Phi_{1}+\Phi^{\dagger}_{2}\Phi_{2}$ $D$
Relevant operators $\Phi^{\dagger}_{3}\Phi_{3}$ $2$
@@
$\Phi^{\dagger}_{1}\Phi_{1}-\Phi^{\dagger}_{1}\Phi_{2}+\Phi^{\dagger}_{2}\Phi_{2}$ $2$
Appendix E Table of scaling dimensions at the isolated fixed points after symmetry-breaking two-loop effects are included
In section IXB, we found that the ring conformal manifold breaks down into isolated fixed points when two-loop fermion field renormalization effects are included. In Tables.11,12, we list the scaling dimensions of the relevant and irrelevant operator(s) at two of these four fixed points. Notice the presence of an operator at the fixed point with its scaling dimension proportional to $a\epsilon^{2}$, second order in $\epsilon$. It suggests that the operator was marginal at the one-loop level.
Table 11: Scaling dimensions of the relevant and irrelevant operators at the isolated fixed point $\hat{V}_{c}=\left(\frac{\epsilon}{2}+a\epsilon^{2}\right)\left(\hat{I}+\sigma_{z}\right)$. Here, $D$ is the spacetime dimensions. The numerical constant $a=\frac{1}{32}$.
@@
Irrelevant operators $-2\Phi^{\dagger}_{2}\Phi_{2}$ $22D-$ 2
Relevant operators $2\Phi^{\dagger}_{1}\Phi_{1}$ $2+2a\epsilon^{2}$
$\Phi^{\dagger}_{1}\Phi_{2}$+ h.c $D-a\epsilon^{2}$
Table 12: Scaling dimensions of the relevant and irrelevant operators at the isolated fixed point $\hat{V}_{c}=\left(\frac{\epsilon}{2}+\frac{a}{2}\epsilon^{2}\right)\left(\hat{I}+\sigma_{x}\right)$. Here, $D$ is the spacetime dimension. The numerical constant $a=\frac{1}{32}$.
@@
Irrelevant operators $-\left(\Phi^{\dagger}_{1}\Phi_{1}+\Phi^{\dagger}_{2}\Phi_{2}\right)+\Phi^{\dagger}_{1}\Phi_{2}$ + h.c $2D-2+a\epsilon^{2}$
$\Phi^{\dagger}_{1}\Phi_{1}-\Phi^{\dagger}_{2}\Phi_{2}$ $D+a\epsilon^{2}$
Relevant operators $\Phi^{\dagger}_{1}\Phi_{1}+\Phi^{\dagger}_{2}\Phi_{2}+\Phi^{\dagger}_{1}\Phi_{2}$ + h.c $2+a\epsilon^{2}$